charges q and 9q separated by distance r how much distance the electric field will be zero form 9q

Grade:12

1 Answers

Kiranchandar
6 years ago
q -------P------------------ 9q
 
Consider two charges q and 9q seperated by a distance ‘r’, also consider a point P located ‘x’ units from 9q where the electric field is zero.
 
Hence at point P,
K*9q/x^2 = K*q/(r-x)^2
 
9/x^2 = 1/(r-x)^2
 
3(r-x) = x
 
x = 3r/4. Hence it must be placed at 3/4th of ‘r’